OpenStudy (antaniece):

If x^2+mx+m is a perfect square trinomial which equation must be true X^2+mx X^2+mx+m(x+2)^2 X^2+mx+m(x+2)^2 X^2+mx+m(x+4)^2

OpenStudy (antaniece):

For the first one it is x^2+mx+m=(x-1)^2

OpenStudy (mathmale):

You might want to practice "completing the square" with a specific problem, and then go back and follow the same steps when only 'm' the variable appears. How would we "complete the square" if given x^2 + 6x? 1) Take half of the coefficient of x, and then square your result: (1/2)(6)=3 => 3^2 = 9 2) Add this square (9) and then subtract this square (9): x^2 + 6x + 9 - 9 3) Rewrite x^2 + 6x + 9 as the square of a binomial: (x+3)^2 4) Rewrite the original expression, x^2 + 6x, as (x+3)^2 - 9 Follow these steps with x^2 + mx + m. Can you find a value for m that makes x^2 + mx + m a "perfect square trinomial?"

OpenStudy (anonymous):

\[\left( \frac{ m }{ 2 } \right)^2=m,m^2=4m,m^2-4m=0,m(m-4)=0,m=0,\] because then it is not a trinomial, then m=4